OpenStudy (anonymous):

the time in paris is 6 hours ahead of new york. how would you adjust your formula if paris time comes out greater than 12? all she gave us was p=v+6 for the first part

OpenStudy (amistre64):

mod your addition

OpenStudy (amistre64):

divide by 12 and take the remainder for your time

OpenStudy (amistre64):

23 oclock = 23/12 = 1 R 11, so 11 oclock

OpenStudy (amistre64):

18 oclock = 18/12 = 1 R 6, soo 6 oclock

OpenStudy (anonymous):

wow thank you very much

OpenStudy (amistre64):

just recall that when R=0 its 12 oclock and not 0 oclock :)

OpenStudy (anonymous):

wait so how would i write that in a formula?

OpenStudy (amistre64):

i forget the proper notation .... i just recall the method

OpenStudy (anonymous):

its ok thanks anyways ill figure out how to write it ina formula

OpenStudy (amistre64):

6/12 = 0 R 6, 6 0clock equation ------- = Q(x) + R(x), time = R(x) 12

OpenStudy (amistre64):

good luck, try looking up stuff like: 5 = 13 (mod 4)

OpenStudy (amistre64):

and modular arithmetic